OpenStudy (anonymous):

x^2/x^2+15 * x+4/x multiply and simplify to lowest terms

OpenStudy (anonymous):

\[\frac{x^2}{x^2+15}\times \frac{x+4}{x}\] the only thing you can do is cancel an \(x\) top and bottom and get \[\frac{x(x+4)}{x^2+15}\]

OpenStudy (anonymous):

so that is the final answer

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

that is the best you can do

OpenStudy (anonymous):

or you can write it as \[\frac{x^2+4x}{x^2+15}\] but they are the same

OpenStudy (anonymous):

thanks can you help me with some more please

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

i'll post a new question in a second greatly apprciated

OpenStudy (anonymous):

are both these answers factored form?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

the first one is in factored form